OpenStudy (anonymous):

A principal of $835 is invested in an account at 3% per quarter simple interest. Which of the following sequences describes the dollar amount of principal in the account at quarterly intervals? A. 66.80, 133.60, 200.40, 267.20, 334.00, ... B. 835.00, 901.80, 968.60, 1,035.40, 1,102.20, ... C. 25.05, 50.10, 75.15, 100.20, 125.25, ... D. 835.00, 860.05, 885.10, 910.15, 935.20, ... E. 835.00, 860.05, 885.70, 911.15, 937.26, ...

OpenStudy (welshfella):

each quarter 0.03 * 835 is added the the principal amount 9 0.03 = 3%) 0.03 * 835 = $25.05

OpenStudy (anonymous):

so c?

OpenStudy (welshfella):

after first quarter principal = 835 + 25.05 = 860.05

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

No c gives the interest you want the principal every quarter

OpenStudy (anonymous):

its either c or d

OpenStudy (welshfella):

which of the options adds 25.05 dollars each quarter?

OpenStudy (anonymous):

its d

OpenStudy (welshfella):

Not c - this gives the interest only

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

thx

OpenStudy (welshfella):

yw